OpenStudy (anonymous):

write the given expression as a monomial containing a single function or a constant. 1-cos^2 phata?

OpenStudy (anonymous):

are you familiar with the trigonometric Pythagorean identities?

OpenStudy (anonymous):

no i do not know where to even begin.

OpenStudy (anonymous):

This is called an identity because it is always true for all values of the variable\[\cos ^{2}\theta+\sin ^{2}\theta=1\]it is called "Pythagorean" since it is derived directly from that famous theorem. Your expression contains two of the three terms in the identity. You just need to rearrange accordingly. try...

OpenStudy (anonymous):

Hint: make one side of the equation I provided look like the one you provided in your question. It is one algebraic manipulation.